# {textblock*}

### v3

@@
Abstract
While graph neural networks (GNNs) operating in hyperbolic spaces have shown promise for modeling hierarchical and complex relational data, a critical limitation often overlooked is their potentially limited discriminative power compared to their Euclidean counterparts or fundamental graph isomorphism tests like the Weisfeiler-Lehman (WL) hierarchy. Ex- isting hyperbolic aggregation schemes, while curvature-aware, may not sufficiently capture the intricate structural differences required to robustly distinguish non-isomorphic graphs owing to non-injective aggregation functions. To address this expressiveness gap in hyper- bolic graph learning, we introduce the Lorentzian Graph Isomorphic Network (LGIN), a novel GNN designed to achieve enhanced discriminative capabilities within the Lorentzian model of hyperbolic space. LGIN proposes a new update rule that effectively combines local neighborhood information with a richer representation of graph structure designed to preserve the Lorentzian metric tensor. This represents a significant step towards building more expressive GNNs in non-Euclidean geometries, overcoming a common bottleneck in current hyperbolic methods. We conduct extensive evaluations across nine diverse benchmark datasets, in- cluding molecular and protein structures. LGIN consistently outperforms or matches state- of-the-art hyperbolic and Euclidean GNNs, showcasing its practical efficacy and validating its superior ability to capture complex graph structures and distinguish between different graphs. To the best of our knowledge, LGIN is the first work to study the framework behind a powerful GNN on the hyperbolic space. The code for our paper can be found at https://github.com/Deceptrax123/LGIN
1 Introduction
Graphs constitute an abstract concept for modeling relationships and structures across diverse scientific and technological domains comprising of social networks(Li et al., 2023), biological systems(Zhang et al., 2021b) and knowledge graphs(Ye et al., 2022). Extracting representational information from graphs by mapping such abstraction to coordinate spaces is critical for downstream tasks such as node classification(Zhao et al., 2021), link prediction(Yun et al., 2021) and graph classification(Xie et al., 2022). Graph Neural Networks ((Kipf & Welling, 2016),(Veličković et al., 2017),(Hamilton et al., 2017)) have achieved notable success in these areas by employing a scheme of feature transformation followed by neighbor aggregation. However, these methods significantly operate in Euclidean spaces, potentially overlooking the hierarchical and multi-relational characteristics inherent to complex graphs.
The limitations of Euclidean space for modeling graphs with hierarchical and multi-scale structures have motivated researchers to explore alternative geometries. Hyperbolic geometry, in particular, has emerged as a promising alternative for graph representation learning. Unlike Euclidean space, hyperbolic space grows exponentially in volume with radius, making it naturally suited for embedding hierarchical data with minimal distortion (Zhu et al., 2023). This property enables hyperbolic embeddings to more accurately capture the intricate relationships present in hierarchical data, such as taxonomies, protein-protein interaction networks, and social hierarchies. Additionally, hyperbolic geometry requires fewer dimensions to represent such structures effectively, resulting in improved computational efficiency and better preservation of structural information (Zhang et al., 2021a).
In parallel, the theoretical framework of Graph Isomorphic Networks (GINs) (Xu et al., 2018) has gained attention due to its ability to capture structural features of graphs and distinguish non-isomorphic graphs. GINs leverage injective aggregation mechanisms and readout functions to ensure that distinct graph structures map to different embeddings, making them as powerful as the Weisfeiler-Lehman (WL) test for graph isomorphism. This has made GINs a popular choice for tasks requiring high expressivity, such as molecular property prediction and community detection. Despite the successes of GINs in Euclidean space, there is a notable absence of such frameworks adapted for hyperbolic spaces, which could potentially unlock even greater representational power by harnessing the unique properties of hyperbolic geometry.
However, extending such frameworks to the hyperbolic space is inherently challenging due to the need to ensure that embeddings remain constrained within the Riemannian manifold. In particular, ensuring that the operations performed during aggregation and transformation preserve the hyperbolic metric is non-trivial. Motivated by the potential to unify high expressive power with geometry suited for hierarchical data, we propose the Lorentzian Graph Isomorphic Network (LGIN). Our design prioritizes maintaining the intrinsic geometric properties and preserving the Lorentzian metric during aggregation and transformation, which presents challenges for achieving strictly injective message-passing. To effectively bridge this expressiveness gap, LGIN introduces a novel node update rule incorporating a cardinality-aware aggregation mechanism. This design is specifically crafted to approximate the strong discriminative capabilities associated with Euclidean Graph Isomorphism Networks (GINs) by preserving crucial information often lost in standard hyperbolic aggregations. Our empirical results demonstrate that LGIN achieves superior discriminative capabilities compared to existing hyperbolic GNNs, including LGCN and other architectures. We further provide a key contribution by empirically demonstrating that the proposed Hyperbolic-Tangent-Hyperbolic approach, when coupled with cardinality-aware aggregation, yields a GNN with expressivity at least approximately equivalent to the 1-dimensional Weisfeiler-Lehman test. By rigorously integrating structural isomorphism testing principles with the inherent geometric inductive biases of hyperbolic space, this work establishes a foundational theoretical framework for designing powerful and highly-discriminative GNNs on non-Euclidean manifolds.
Therefore, we summarize our key contributions as follows:
@@ 1 Introduction
  • We propose the theoretical framework for Lorentzian Graph Isomorphic Netork(LGIN), with a modified cardinality-scaled aggregation rule to preserve the cardinality information of multisets and a new update rule, as an approximately powerful alternative to Graph Isomorphic Network(GIN) and other state of the art hyperbolic graoph neural networks such as LGCN.
@@
2 Related Work
In this section we describe the related work with respect to Hyperbolic GNNs, mathematical frameworks explaining powerful GNNs and applications of hyperbolic graph neural networks.
2.1 Advances in Hyperbolic Graph Neural Networks
Hyperbolic graph neural networks, as discussed above, have emerged as a powerful method for modeling complex and hierarchical structures. (Liu et al., 2019) generalized graph neural networks to be manifold-agnostic and demonstrated that hyperbolic graph neural networks are more efficient at capturing structural properties than their Euclidean counterpart. (Khatir et al., 2022) proposed the Poincare disk model as the search space to apply all approximations on the disk, thus eliminating the need for inter-space transformations. This framework introduces a hyperbolic normalization layer that simplifies the entire hyperbolic model to a Euclidean model cascaded with the normalization layer, maintaining the advantages of both geometric approaches. (Fu et al., 2023) introduced ACE-HGNN which dynamically learns the optimal curvature based on the input graph and downstream tasks. Using a multi-agent reinforcement learning framework with two agents—ACE-Agent and HGNN-Agent—for learning curvature and node representations respectively, this model adapts to the specific hierarchical structures present in different graphs rather than using a manually fixed curvature value. (Zhang et al., 2021a) introduced Lorentzian Graph Convolutional Networks(LGCN) for modeling hierarchical architectures. They design a neighborhood aggregation method based on the centroid of Lorentzian distance to constrain embeddings within the hyperboloid. Further, they theoretically prove that some proposed graph operations are equivalent to those defined in other hyperbolic models such as the Poincare ball model. (Choudhary et al., 2023) introduced H-GRAM, a meta-learning framework specifically designed for hyperbolic graph neural networks. The model learns transferable information from local subgraphs in the form of hyperbolic meta gradients and label hyperbolic protonets thereby enabling faster learning over new tasks dealing with disjoint subgraphs, thus addressing the generalization challenge.
2.2 Applications of Hyperbolic Graph Neural Networks
(Ramirez et al., 2025) used a Fully Hyperbolic Graph Neural Network(FHGNN) to embed functional brain connectivity graphs derived from magnetoencephalography (MEG) data into low dimensions on a Lorentz model of hyperbolic space. (Wu et al., 2021) developed a graph-based Quantitative Structure-Activity Relationship(QSAR) method by building a hyperbolic relational graph convolution network. It leverages both molecular structure and molecular descriptors to achieve state-of-the-art performance on 11 drug discovery-related datasets. (Bao et al., 2025) introduced Hyperbolic Graph Convolutional Network-based Knowledge Graph Embedding (HGCGE) which addresses challenges in traditional GCN-based KGE methods, such as oversmoothing and high distortion in Euclidean space. It employs GCN operations in hyperbolic space with Möbius transformations to embed entities and relationships in the Poincaré model, enhancing hierarchical data representation. The proposed scoring function improves entity distinction across relationships, achieving superior performance on multiple datasets, even at low dimensions and training rounds.
2.3 Graph Expressiveness
(Xu et al., 2018) presented a theoretical framework for analyzing the expressive power of GNNs. Their analysis demonstrated that popular GNN variants such as Graph Convolutional Networks and GraphSAGE cannot distinguish certain simple graph structures, limiting their discriminative capabilities. The researchers then developed a simple architecture—the Graph Isomorphism Network—that is provably the most expressive among the class of message-passing GNNs and as powerful as the 1-dimensional Weisfeiler-Lehman(WL) graph isomorphism test. (Hevapathige & Wang, 2023) showed that partitioning a graph into sub-graphs that preserve structural properties provides a powerful means to exploit interactions among different structural components of the graph. The researchers proposed Graph Partitioning Neural Networks(GPNNs) that combines structural interactions via permutation invariant graph partitioning to enhance graph representation learning.
3 Preliminaries
In this section, we formally define the key terms and establish the theoretical framework underlying our proposed work. For a deeper understanding of geometric topology, refer to (Martelli, 2016).
3.1 Graph Preliminaries
Weisfeiler-Lehman(WL) Test. The Weisfeiler-Lehman (WL) test is a graph isomorphism heuristic that iteratively refines node representations based on local neighborhood aggregation. Given a graph, the test assigns initial labels to nodes and updates them iteratively by incorporating labels from neighboring nodes. This process continues until convergence, producing a unique representation for each node that captures its structural role within the graph(Morris et al., 2019). The WL test is widely used to distinguish non-isomorphic graphs efficiently and serves as a foundation for many graph neural networks (GNNs)(Xu et al., 2018).
3.2 Definitions of Key Topological Concepts
Definition 3.1 (Differential Manifold).
Definition 3.2 (Tangent Space).
Definition 3.3 (Metric Tensor).
$g:T\mathcal{M}\times T\mathcal{M}\to\mathbb{R}$
$g_{p}(X,Y)=g_{p}(Y,X),\quad\forall X,Y\in T_{p}\mathcal{M},$
Definition 3.4 (Curvature).
$R(X,Y)Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z-\nabla_{[X,Y]}Z,$ (1)
$K(X,Y)=\frac{\langle R(X,Y)Y,X\rangle_{g}}{\|X\|_{g}^{2}\|Y\|_{g}^{2}-\langle X,Y\rangle_{g}^{2}}.$ (2)
$R(X,Y)Z=K\left(\langle Y,Z\rangle_{g}X-\langle X,Z\rangle_{g}Y\right).$ (3)
Definition 3.5 (Isometry).
$\langle v,w\rangle=\langle df_{p}(v),df_{p}(w)\rangle$ (4)
Definition 3.6 (Parallel Transport).
3.3 Hyperboloid Model
$\langle x,y\rangle_{\eta}=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$ (5)
$\mathbb{H}^{n}=\left\{x\in\mathbb{R}^{n+1}\mid\langle x,x\rangle_{\eta}=-1,x_{0}>0\right\}$ (6)
$\exp_{p}:T_{p}M\longrightarrow M$
is defined as follows: A vector $v\in T_{p}\mathcal{M}$ determines a maximal geodesic $\gamma_{v}:\mathbb{R}\to\mathcal{M}$ with $\gamma_{v}(0)=p$ and $\gamma_{v}^{\prime}(0)=v$. We set $\exp_{p}(v)=\gamma_{v}(1)$. The logarithmic map $\log_{p}$ is the inverse of the exponential map. In the hyperboloid space, the exponential and logarithmic maps are given by equations 7 and 8 respectively.
$\exp_{p}(v)=\cosh\left(\sqrt{|c|}\|v\|_{\eta}\right)p+\sinh\left(\sqrt{|c|}\|v\|_{\eta}\right)\frac{v}{\|v\|_{\eta}}$ (7)
$\log_{p}(q)=\frac{d_{\mathbb{H}}(p,q)}{\sinh\left(\sqrt{|c|}d_{\mathbb{H}}(p,q)\right)}\left(q-\cosh\left(\sqrt{|c|}d_{\mathbb{H}}(p,q)\right)p\right)$ (8)
where
$d_{\mathbb{H}}(p,q)=\frac{1}{\sqrt{|c|}}\cosh^{-1}(-\langle p,q\rangle_{\eta})$ (9)
$PT^{c}_{p\rightarrow q}(v)=v-\frac{c\langle q,v\rangle_{\eta}}{1+c\langle p,q\rangle_{\eta}}(p+q)$ (10)
4 Proposed Framework
@@ 4.1 Overview
$f^{*}g=g\quad\text{and}\quad f\circ\Phi(G_{1})=\Phi(G_{2})$ (13)
@@
  • A powerful graph neural networks requires strictly injective aggregation and readout functions(Xu et al., 2018). However, directly translating strictly injective operations to Riemannian manifolds while preserving geometric properties is non-trivial. Our goal is to design a network that achieves high expressive power by carefully approximating these properties within the Lorentzian geometry.
@@
4.2 Injectivity of Lorentz Transformation
We show the injectivity of the Lorentz Transformation on the manifold and through moment injectivity proposed by (Amir et al., 2023). To ensure the injectivity of the Lorentz transformation, we follow the hyperbolic-tangential-hyperbolic approach (Zhang et al., 2021a). This is because the tangent space at a point is locally isometric to Euclidean space. The Lorentz transformation used in our proposed method is given by:
$\mathbf{W}\otimes_{c}^{\mathcal{L}}\mathbf{x}^{\mathcal{L}}:=\exp_{o}^{c}\left(0,\mathbf{W}\log_{o}^{c}\left(\mathbf{x}^{\mathcal{L}}\right)_{[1:n]}\right),$ (14)
@@ 4.2.1 Injectivity on the Manifold
4.2.2 Moment Injectivity
From (Amir et al., 2023), we have the following result for the Euclidean space: Consider shallow neural networks $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ of the form
$f(\mathbf{x};A,\mathbf{b})=\sigma(A\mathbf{x}+\mathbf{b}),\quad A\in\mathbb{R}^{m\times d},\ \mathbf{b}\in\mathbb{R}^{m},$ (15)
$\hat{f}:\mathcal{M}_{\leq n}(\Omega)\rightarrow\mathbb{R}^{m}$
defined by
[…]
Therefore,we use a shallow(2-Layer) Lorentzian network based on transformation equation 14 with a pointwise Tanh/Sigmoid function and ensuring $\mathbf{W}$ has a full column rank.
4.3 Lorentzian Graph Isomorphic Network
Based on the conditions mentioned in section 4.1, proof in section 4.2 and corollary 6 in (Xu et al., 2018), we finally define the update rule for the Lorentzian Graph Isomorphic Network(LGIN) as follows:
$\mathbf{x}^{\mathcal{L}^{k}}=\text{LT}^{k}(\exp_{o}^{c}((1+\epsilon^{k}).\mathcal{P}_{x^{{L}^{k-1}}\longrightarrow T}(\log_{o}^{c}(\mathbf{x}^{\mathcal{L}^{k-1}}))+\log_{o}^{c}(T(\mathbf{x}^{\mathcal{L}^{k-1}}))))$ (18)
$T(\mathbf{x}_{i}^{\mathcal{L}^{k-1}}):=\frac{\sqrt{c}\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{x}_{j}^{\mathcal{L}}}{(1+|\mathcal{N}_{i}|)\lVert\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{x}_{j}^{\mathcal{L}}\rVert_{\mathcal{L}}}$ (19)
$\alpha_{ij}=\text{softmax}_{j\in\mathcal{N}(i)}\left(-d_{c}^{2}(\mathbf{x}_{i}^{\mathcal{L}},\mathbf{x}_{j}^{\mathcal{L}})\right)$ (20)
4.3.1 Justification of Lorentz Centroid
Equation 19 ensures embeddings reside on the hyperbolic manifold after aggregation i.e the preservation of metric tensors. Normalization prevents the node embeddings from collapsing into the origin or drifting too far. However, since attention-based aggregators are ot strictly injective, it may lead to non-isomorphic graphs having the same embeddings. To mitigate this, we modify the Lorentz aggregation function to preserve the cardinality information of multisets during aggregation which can approximate the injective property(Zhang & Xie, 2020).
4.3.2 Justification of Parallel Transport
5 Experiments
5.1 Experimental conditions
5.1.1 Datasets
For our experiments, we use 6 Bioinformatics datasets and 3 MoleculeNet datasets. The bioinformatics datasets include, namely, Proteins, Mutag, NCI1, PTC, Enzymes, and DD(Yanardag & Vishwanathan, 2015). The MoleculeNet datasets include, namely, BBBP, BACE and HIV(Wu et al., 2018). The initial features of the Bioinformatics datasets are mostly the node degree, while for MoleculeNet, each node is assigned 9 features such as atomic number, chirality, formal charge, or whether the atom is in a ring or not.
5.1.2 Hyperparameters and model configuration
In our proposed method, we evaluate on sets of variants. One with a constant curvature $c$ and the other with variable curvature, which is essentially a trainable parameter. All initial curvatures were set to 4. The trainable parameter $\epsilon$ was initially set to 0.1. We use the Riemannian Adam(Bécigneul & Ganea, 2018) to train our model with a cosine annealing learning rate scheduler with warm restarts. We use the same model architecture for all datasets, which is a 3-layer Lorentzian Graph Isomorphic Network predicting 128, 256 and 512 features respectively. To ensure stability of gradient flow, we make use of gradient clip after gradient computation at each step. The number of Lorentz Transformation layers have been set to 2.
5.2 Evaluations
In this section, we compare our framework against several baselines as mentioned in the previous section. We mainly compare our model with Graph Isomorphic Network(GIN), a powerful GNN framework defined for the Euclidean space and its variants. Unless otherwise mentioned, for Bioinformatics datasets, we record the mean and standard deviation of the test accuracy on 10 splits. For MoleculeNet datasets, we consider the average accuracy on 3 splits.
5.2.1 Evaluations against Graph Isomorphic Network
Table 1: Performance of GNN Variants against LGIN. The best score for all variants are marked in bold along with LGINs comparable with GIN variants based on a paired t-test with significance level of 10%. Models performing significantly better are marked in bold with an asterisk. Evaluation standard defined as per (Xu et al., 2018)
@@ 5.2.1 Evaluations against Graph Isomorphic Network
Table 1 contains the performance of our model against powerful baselines. The performance of various GNN variants against LGIN was evaluated across four datasets: MUTAG, PROTEINS, PTC, and NCI1.The proposed LGIN variants demonstrate competitive performance across diverse datasets. Notably, LGIN with Fixed Curvature achieves the highest accuracy on the MUTAG dataset (90.6 $\pm$ 3.9) and the PTC dataset (68.4 $\pm$ 5.2), outperforming several existing models. Furthermore, LGIN with Variable Curvature attains the best performance on the PROTEINS dataset (79.1 $\pm$ 4.4), showing statistically significant improvement ($p<0.1$). In the case of the NCI1 dataset, the Weisfeiler-Lehman (WL) subtree kernel achieved the best performance with 86.0 $\pm$ 1.8, while the PATCHYSAN model showed competitive performance on the MUTAG dataset with 92.6 $\pm$ 4.2. Across other datasets, several GIN variants such as SUM-MLP (GIN-$\varepsilon$) and SUM-1-LAYER also reported consistent results. The performance gains observed in LGIN, particularly in datasets with varying node sizes and graph complexities, underscore its ability to effectively model hierarchical relationships by leveraging Lorentzian geometry. Overall, the results demonstrate that incorporating curvature into Graph Isomorphism Networks enhances representational capacity, enabling more effective learning across diverse graph structures.
5.2.2 Performance on Enzymes, Proteins and DD against state of the art graph neural networks
Table 2: Detailed Analysis on Proteins, Enzymes and DD. The best accuracy on the test splits are marked in bold while the second best score is underlined. Hyphen indicates the scores were not reported by the authors. Hyperbolic GNN baselines have been marked with an asterisk(*).
Model $L$ Test Acc. $\pm$ s.d. Train Acc. $\pm$ s.d. #Epoch

### v4

# Lorentzian Graph Isomorphic Network

@@
Abstract.
While hyperbolic GNNs show promise for hierarchical data, they often have limited discriminative power compared to Euclidean counterparts or the WL test, due to non-injective aggregation. To address this expressivity gap, we propose the Lorentzian Graph Isomorphic Network (LGIN), a novel HGNN designed for enhanced discrimination within the Lorentzian model. LGIN introduces a new update rule that preserves the Lorentzian metric while effectively capturing richer structural information. This marks a significant step towards more expressive GNNs on Riemannian manifolds. Extensive evaluations across nine benchmark datasets demonstrate LGIN’s superior performance, consistently outperforming or matching state-of-the-art hyperbolic and Euclidean baselines, showcasing its ability to capture complex graph structures. LGIN is the first to adapt principles of powerful, highly discriminative GNN architectures to a Riemannian manifold. The code for our paper can be found at https://github.com/Deceptrax123/LGIN
Riemannian Manifolds, Graph Neural Networks, Graph Isomorphism
conference: International Conference on Management of Data; June 22–27, 2025; Berlin, Germanyjournalyear: 2025ccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning by classification
1. Introduction
Graphs model complex relationships, but standard Euclidean Graph Neural Networks (GNNs) ((kipf2016semi, ),(velivckovic2017graph, )) struggle with graphs exhibiting strong hierarchical or multi-relational structures. Hyperbolic geometry, with its exponential volume growth, is better suited for embedding such data with minimal distortion (zhu2023lorentzian, ). In parallel, Graph Isomorphic Networks (GINs) (xu2018powerful, ) achieve high expressivity in Euclidean space, equivalent to the 1-dimensional Weisfeiler-Lehman test, through injective aggregation. However, extending GIN’s expressivity to hyperbolic space is challenging due to the need to preserve geometric properties and metric constraints during message passing.
Motivated to bridge this gap, we propose the Lorentzian Graph Isomorphic Network (LGIN), the first attempt at an isomorphic GNN incorporating constant curvature. LGIN employs a novel hyperbolic-tangent-hyperbolic approach to address metric preservation. To approximate the discriminative power of Euclidean GINs, LGIN introduces a modified cardinality-scaled aggregation rule, preserving multiset information often lost in hyperbolic settings. This design yields a GNN with expressivity at least approximately equivalent to the 1-dimensional Weisfeiler-Lehman test. Empirical results demonstrate LGIN’s superior discriminative capabilities and performance compared to existing hyperbolic GNNs like LGCN across various benchmarks. Therefore, we summarize our key contributions as follows:
@@ 1. Introduction
  • We propose the theoretical framework for Lorentzian Graph Isomorphic Netork(LGIN), with a modified cardinality-scaled aggregation rule to preserve the cardinality information of multisets and a new update rule, as an approximately powerful alternative to Graph Isomorphic Network(GIN) and other state of the art hyperbolic graph neural networks such as LGCN.
@@
2. Related Work
In this section, we outline related work concerning Hyperbolic Graph Neural Networks (HGNNs) and Graph Expressiveness. Hyperbolic Graph Neural Networks have shown promise for modeling complex and hierarchical structures, benefiting from hyperbolic space’s natural ability to embed such data with low distortion (liu2019hyperbolic, ). Various approaches have been explored, including manifold-agnostic generalizations (liu2019hyperbolic, ), utilizing specific models like the Poincare disk with normalization layers (khatir2022unification, ), adapting curvature based on data (fu2023adaptive, ), and developing methods like Lorentzian Graph Convolutional Networks (LGCN) which use centroid aggregation in the Lorentz model (zhang2021lorentzian, ). HGNNs have found successful applications in diverse domains, including brain connectivity, drug discovery, and knowledge graphs.
Understanding the expressive power of GNNs is crucial for distinguishing graph structures. The 1-dimensional Weisfeiler-Leman (WL) test provides a benchmark for graph isomorphism testing. Graph Isomorphism Networks (GINs) (xu2018powerful, ) represent a class of message-passing GNNs in Euclidean space that are provably as powerful as the 1-WL test, primarily due to their use of injective aggregation functions. This highlights the importance of aggregation mechanisms for achieving high discriminative capabilities. We have elaborated on the related work in appendix F.
3. Proposed Framework
@@ 3.1. Overview
(3) $f^{*}g=g\quad\text{and}\quad f\circ\Phi(G_{1})=\Phi(G_{2})$
@@
  • A powerful graph neural networks requires strictly injective aggregation and readout functions(xu2018powerful, ). However, directly translating strictly injective operations to Riemannian manifolds while preserving geometric properties is non-trivial. Our goal is to design a network that achieves high expressive power by carefully approximating these properties within the Lorentzian geometry.
@@
3.2. Injectivity of Lorentz Transformation
We show the injectivity of the Lorentz Transformation on the manifold and through moment injectivity proposed by (amir2023neural, ). To ensure the injectivity of the Lorentz transformation, we follow the hyperbolic-tangential-hyperbolic approach (zhang2021lorentzian, ). This is because the tangent space at a point is locally isometric to Euclidean space. The Lorentz transformation used in our proposed method is given by:
(4) $\mathbf{W}\otimes_{c}^{\mathcal{L}}\mathbf{x}^{\mathcal{L}}:=\exp_{o}^{c}\left(0,\mathbf{W}\log_{o}^{c}\left(\mathbf{x}^{\mathcal{L}}\right)_{[1:n]}\right),$
@@ 3.2.1. Injectivity on the Manifold
Table 1. Performance of GNN Variants against LGIN. The best score for all variants are marked in bold along with LGINs comparable with GIN variants based on a paired t-test with significance level of 10%. Models performing significantly better are marked in bold with an asterisk. Evaluation standard defined as per (xu2018powerful, )
@@
3.2.2. Moment Injectivity
From (amir2023neural, ), we have the following result for the Euclidean space: Consider shallow neural networks $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ of the form
(5) $f(\mathbf{x};A,\mathbf{b})=\sigma(A\mathbf{x}+\mathbf{b}),\quad A\in\mathbb{R}^{m\times d},\ \mathbf{b}\in\mathbb{R}^{m},$
$\hat{f}:\mathcal{M}_{\leq n}(\Omega)\rightarrow\mathbb{R}^{m}$
defined by
[…]
where $x^{\mathcal{L}}$ resides in the tangent space. Therefore,we use a shallow(2-Layer) Lorentzian network based on transformation equation 4 with a pointwise Tanh/Sigmoid function and ensuring $\mathbf{W}$ has a full column rank.
3.3. Lorentzian Graph Isomorphic Network
Based on the conditions mentioned in section 3.1, proof in section 3.2 and corollary 6 in (xu2018powerful, ), we finally define the update rule for the Lorentzian Graph Isomorphic Network(LGIN) as follows:
(8) $\mathbf{x}^{\mathcal{L}^{k}}=\text{LT}^{k}(\exp_{o}^{c}((1+\epsilon^{k}).\mathcal{P}_{x^{{L}^{k-1}}\longrightarrow T}(\log_{o}^{c}(\mathbf{x}^{\mathcal{L}^{k-1}}))+\log_{o}^{c}(T(\mathbf{x}^{\mathcal{L}^{k-1}}))))$
(9) $T(\mathbf{x}_{i}^{\mathcal{L}^{k-1}}):=\frac{\sqrt{c}\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{x}_{j}^{\mathcal{L}}}{(1+|\mathcal{N}_{i}|)\lVert\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{x}_{j}^{\mathcal{L}}\rVert_{\mathcal{L}}}$
(10) $\alpha_{ij}=\text{softmax}_{j\in\mathcal{N}(i)}\left(-d_{c}^{2}(\mathbf{x}_{i}^{\mathcal{L}},\mathbf{x}_{j}^{\mathcal{L}})\right)$
3.3.1. Justification of Lorentz Centroid
Equation 9 ensures embeddings reside on the hyperbolic manifold after aggregation i.e the preservation of metric tensors. Normalization prevents the node embeddings from collapsing into the origin or drifting too far. However, since attention-based aggregators are ot strictly injective, it may lead to non-isomorphic graphs having the same embeddings. To mitigate this, we modify the Lorentz aggregation function to preserve the cardinality information of multisets during aggregation which can approximate the injective property(zhang2020improving, ).
3.3.2. Justification of Parallel Transport
4. Experiments
4.1. Evaluations
4.1.1. Evaluations against Graph Isomorphic Network
Table 2. Detailed Analysis on Proteins, Enzymes and DD. Test accuracy $\pm$ standard deviation. The best score is in bold and the second best is underlined. Hyphen indicates scores not reported. Hyperbolic GNN baselines are marked with an asterisk(*).
Model $L$ Test Acc. $\pm$ s.d. Train Acc. $\pm$ s.d.